Determine the roots of √2x2 +3x+√2=0 using the method of completing squares
Answer by stanbon(75887)

Determine the roots of
√2x^2 +3x+√2 = 0 using the method of completing squares
----
Divide thru by sqrt(2) to get:
x^2 + (3sqrt(2)/2)x = -1
---
Complete the square:
x^2 +(3sqrt(2)/2)x + (3sqrt(2)/4)^2 = -1+(3sqrt(2)/4)^2
---
(x+3sqrt(2)/4)^2 = -1+(9/8)
---
(x+3sqrt(2)/4)^2 = 1/8
---
x = (-3/4)sqrt2+(1/8) or x = (3/4)sqrt(2)+(1/8)
================================================
Cheers,
stan H.